OpenStudy (anonymous):

f(x) 2sqrt{x} cutoff

OpenStudy (hba):

What ? \[f(x)=2\sqrt{2}\] What do you have to find ?

OpenStudy (anonymous):

cutoff

OpenStudy (anonymous):

sry!! what do u mean by (CutOFF) ?!!!

OpenStudy (anonymous):

derived...

OpenStudy (anonymous):

\[2\sqrt{x}\]

OpenStudy (anonymous):

hmmmm.... look f(x)=2sqrtx ........................ since 2sqrtx=2x^1/2 Therefore: f*(x)= (1/2.2).X^(1-1/2) Therefore f*(x)=X^-1/2=1/sqrtx

OpenStudy (anonymous):

\[2\sqrt{x} = \frac{ 1 }{ 2\sqrt{x} }\] ic correct

OpenStudy (anonymous):

No there is no 2 it will be \[\frac{1 }{ \sqrt{x} }\] only

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did u gt it or u want me to explain more ?!!

OpenStudy (anonymous):

no i got it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

can u help me one question

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

wait one minute